OpenStudy (anonymous):

complete perfect square 2z^2+24z

OpenStudy (anonymous):

Is that one side?

OpenStudy (anonymous):

yea there is more too it but is irrelevant i cant remember how too

OpenStudy (anonymous):

Are you trying to find the perimeter

OpenStudy (anonymous):

the whole problem is 2x^2+2y^2+2z^2=8x-24z+1

OpenStudy (anonymous):

question is show the equation that represents a sphere and find its center and radius

OpenStudy (mertsj):

2z^2+24z=2(z^2+12z) 2(z^2+12z+36)=2(z+6)^2

OpenStudy (anonymous):

yes where did u get the 36

OpenStudy (anonymous):

from that was problem

OpenStudy (anonymous):

thats no help wolfram could have told me that